In three days, Ana traveled 146 miles down the Mississippi River in her kayak with 30 hours of paddling. The first day, she averaged 6mph, the second day 5mph, and the third day 4mph. If her distance on the third day was equal to her distance on the first day, then how many hours did she paddle each day?
Found 2 solutions by josgarithmetic, mananth:
Answer by josgarithmetic(39616)

Uniform Rates for travel uses the basic rule RT=D rate time distance.

________________speed_________time_________distance
Day ONE__________6_____________t___________(____)
Day TWO__________5___________(____)________(____)
Day Three________4_____________t___________(____)
TOTALS________________________30___________146

Fill the spots in the data table and form the two needed equations.

Answer by mananth(16946)

let her distance on the third day and distance on the first day,=x miles
Second day she travels (146-2x)

The first day, she averaged 6mph,
time first day = x/6
the second day 5mph,
time second day = (146-2x)/5
and the third day 4mph.
time third day = x/4
total time = 30 hours
x/6 + (146-2x)/5 + x/4 = 30
multiply equation by LCM 60
10x+12(146-2x)+15x= 1800
10x+1752-24x+15x=1800
x=48
x=48
first &third day 48 miles
balance on second day

then how many hours did she paddle each day?